OpenStudy (anonymous):

Nancy has a sewing project which calls for panels which use 1 1/4 yd of fabric each. She has 4 1/2 yd of fabric. How many panels can she cut from the fabric? A. 2 B. 3 C. 4 D. 5

OpenStudy (anonymous):

c

OpenStudy (anonymous):

The number of 1 1/4 yd panels that are in 4 1/2 yards of fabric is (4 1/2)/(1 1/4) = (9/2)/(5/4) = (9/2)*(4/5) = 18/5= 3.6. Therefore, she can but 3 panels full from the fabric.